OpenStudy (anonymous):

Simplify your answer (3x+5)(7x^2+7x+4)

OpenStudy (anonymous):

\[(3x+5)(7x^2+7x+4)=\] \[(21x^3+21x^2+12x+35x^2+35x+20)=\] \[21x^3+21x^2+35x^2+12x+35x+20=\] \[21x^3+56x^2+47x+20\]

OpenStudy (anonymous):

your answer is wrong...it says to simplify...what is the answer?????????????